# Optimization-Free Fast Optimal Control: Bang-Ride Property, Monotonicity, and Applications to Fast Battery Charging

### v1

@@ 7.2 Proof of Proposition 3
Assumption 3.4b leads to continuous $\lambda$ in $(0,t_{f})$, e.g., see [14, Thm. 4.2] and [13, Eq. (71) in Chap. 6]. Moreover, Assumption 3.4a shows that $\lambda$ is right continuous at 0 and left continuous at $t_{f}$. Therefore, $\lambda$ is continuous at any $t\in[0,t_{f}]$, leading to $\eta(\tau_{i})=0$ for any $i$. Besides, recall that the effect of $h^{\star}_{x}$ on $\lambda(t_{f}^{-})$ in (9a) is caused by the switching of state constraints at $t_{f}$, following the analysis in [31, Appendix B]. Then Assumption 3.4 also implies $\gamma=0$ in (9a).
@@ 7.3 Proof of Theorem 2
Then Assumptions 4.2 and 4.3 show that $\Gamma^{\star}[t]$ is Metzler and $\lambda_{0}l_{x}^{\star}[t]\geqslant 0$ for a.e. $t\in[0,t_{f}]$. Then, based on [15, Lemma VIII.1], the differential equation of $\zeta$ is a positive system, i.e., $\zeta(0)\geqslant 0$ leads to $\zeta(s)\geqslant 0$ for a.e. $s\in[0,t_{f}]$, and thus $\lambda(t)\geqslant 0$ for a.e. $t\in[0,t_{f}]$. Also note that $\lambda(t)\not=0$ for a.e. $t$ under Assumption 3.3, according to Proposition 3. Combining the above with the assumption $g^{\star}[t]>0$ proves the result.

### v3

@@ 7.2 Proof of Proposition 3
Assumption 3.4b leads to continuous $\lambda$ in $(0,t_{f})$, e.g., see [15, Thm. 4.2] and [14, Eq. (71) in Chap. 6]. Moreover, Assumption 3.4a shows that $\lambda$ is right continuous at 0 and left continuous at $t_{f}$. Therefore, $\lambda$ is continuous at any $t\in[0,t_{f}]$, leading to $\eta(\tau_{i})=0$ for any $i$. Besides, recall that the effect of $h^{\star}_{x}$ on $\lambda(t_{f}^{-})$ in (9a) is caused by the switching of state constraints at $t_{f}$, following the analysis in [32, Appendix B]. Then Assumption 3.4 also implies $\gamma=0$ in (9a).
@@ 7.3 Proof of Theorem 2
Then Assumptions 4.2 and 4.3 show that $\Gamma^{\star}[t]$ is Metzler and $\lambda_{0}l_{x}^{\star}[t]\geqslant 0$ for a.e. $t\in[0,t_{f}]$. Then, based on [16, Lemma VIII.1], the differential equation of $\zeta$ is a positive system, i.e., $\zeta(0)\geqslant 0$ leads to $\zeta(s)\geqslant 0$ for a.e. $s\in[0,t_{f}]$, and thus $\lambda(t)\geqslant 0$ for a.e. $t\in[0,t_{f}]$. Also note that $\lambda(t)\not=0$ for a.e. $t$ under Assumption 3.3, according to Proposition 3. Combining the above with the assumption $g^{\star}[t]>0$ proves the result.